# A High-Quality Dataset and Reliable Evaluation for Interleaved Image-Text Generation

### v1

@@ E.2 Interleaved Image-Text Answer Evaluation Dimensions
Text Response Quality (0-5 points): This dimension only focuses on the correspondence between the text response and the question, whether the content precisely matches the user’s needs, and whether the information is complete and error-free. It does not consider the output of any other dimensions and evaluation criteria.
@@
Image Content Quality (0-5 points): This dimension only focuses on the correspondence between the image content and the question (considering the content of the picture, the degree to which the image content answers the question). Whether the key parts are retained, and whether there is an obvious lack of objects.
@@
Image Aesthetic Quality (0-5 points): This dimension only focuses on the performance of the basic generation technology of the image (do not consider the content of the picture). Whether it is clear, whether there are blurred, noisy or out-of-focus areas, truncations or damages (that is, the judgment of image aesthetics and subjective quality).
@@
Text-Image Synergy (0–5 points): This dimension evaluates the degree of alignment and complementarity between the textual and visual components of a response. It focuses not only on how well the entities or scenes described in the text are accurately and completely depicted in the image, but also on whether the text and image together form a coherent and mutually supportive answer to the question.
@@
Appendix F All Prompts Used in this Work
F.1 Prompts Used in SEIR Method
@@
F.2 Interleaved Image-text Answer Evaluation Prompt Used by MLLM
@@
[Text Response Quality: *; Image Response Quality: *; Image Aesthetic Quality: *; Text-Image Synergy: *]
@@
Appendix G Benchmark Samples
@@
Appendix H Human Annotation Platform
@@
Appendix I Limitations of This Study
While our work introduces InterSyn, the first large-scale, instruction-following dataset for multi-turn, interleaved image-text dialogues, and proposes SynJudge, a comprehensive automatic evaluator emphasizing image-text synergy, several limitations remain that suggest directions for future improvement.

### v2

@@ F.2 Interleaved Image-Text Answer Evaluation Dimensions
Text Content Completeness (0-5 points): This dimension only focuses on the correspondence between the text response and the question, whether the content precisely matches the user’s needs, and whether the information is complete and error-free. It does not consider the output of any other dimensions and evaluation criteria.
@@
Image Content Completeness (0-5 points): This dimension only focuses on the correspondence between the image content and the question (considering the content of the picture, the degree to which the image content answers the question). Whether the key parts are retained, and whether there is an obvious lack of objects.
@@
Image Quality (0-5 points): This dimension only focuses on the performance of the basic generation technology of the image (do not consider the content of the picture). Whether it is clear, whether there are blurred, noisy or out-of-focus areas, truncations or damages (that is, the judgment of image aesthetics and subjective quality).
@@
Image-Text Synergy (0–5 points): This dimension evaluates the degree of alignment and complementarity between the textual and visual components of a response. It focuses not only on how well the entities or scenes described in the text are accurately and completely depicted in the image, but also on whether the text and image together form a coherent and mutually supportive answer to the question.
@@
Appendix G All Prompts Used in this Work
G.1 Prompts Used in SEIR Method
@@
G.2 Interleaved Image-text Answer Evaluation Prompt Used by MLLM
@@
[Text Content Completeness: *; Image Content Completeness: *; Image Quality: *; Image-Text Synergy: *]
@@
Appendix H Benchmark Samples
@@
Appendix I Human Annotation Platform
@@
Appendix J Limitations of This Study
While our work introduces InterSyn, the first large scale, instruction-following dataset for multi-turn, interleaved image-text dialogues, and proposes SynJudge, a comprehensive automatic evaluator emphasizing image-text synergy, several limitations remain that suggest directions for future improvement.